Healthy Harry’s Juice Bar has the following cost schedules:

Q (VATS)

VARIABLE COST

TOTAL COST


$ 0

$ 30

1

10

40

2

25

55

3

45

75

4

70

100

5

100

130

6

135

165

a. Calculate average variable cost, average total cost, and marginal cost for each quantity.

b. Graph all three curves. What is the relationship between the marginal-cost curve and the average total- cost curve? Between the marginal-cost curve and the average-variable-cost curve? Explain.


1
Expert's answer
2022-02-28T14:01:57-0500

a). AVC, TAC and MC

Total variable cost=0+10+25+45+70+100+135=385

Total cost=30+40+55+75+100+130+165=595

Total units=0+1+2+3+4+5+6=21


"Average variable cost=\frac{Total variable cost}{units}=\frac{385}{21}=18.33"


"Average Total cost=\frac{Total cost}{units}=\frac{595}{21}=28.33"


Marginal cost:

1st additional unit; 40-30=10

2nd additional unit; 55-40=15

3rd additional unit; 75-55=20

4th additional unit; 100-75=25

5th additional unit; 130-100=30

6th additional unit; 165-130=35


b). Graphical illustration